OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 3x + 2y = 7 4x - 3y = -2 Will give Medals!!!

OpenStudy (anonymous):

are you replying?

OpenStudy (texaschic101):

I, myself, would use elimination just because if you used substitution, you would be dealing with fractions 3x + 2y = 7 -->(3) 4x - 3y = -2 -->(2) --------------- 9x + 6y = 21 (result of multiplying by 3) 8x - 6y = -4 (result of multiplying by 2) ---------------add 17x = 17 x = 1 now sub 1 in for x in either equation to find y 3x + 2y = 7 3(1) + 2y = 7 3 + 2y = 7 2y = 7 - 3 2y = 4 y = 2 check.. 4x - 3y = -2 4(1) - 3(2) = -2 4 - 6 = -2 -2 = -2 (correct) solution is : (1,2)